# Distributionally Robust Markov Games with Average Reward

### v2

@@ Abstract
We study distributionally robust Markov games (DR‑MGs) with the average‑reward criterion, a crucial framework for multi-agent decision-making under uncertainty over extended horizons. We first establish a connection between the best-response policies and the optimal policies for the induced single‑agent problems. Under a standard irreducible assumption, we derive a correspondence between the optimal policies and the solutions of the robust Bellman equation, and derive the existence of stationary Nash Equilibrium (NE) based on these results. We also study a more general weakly communicating setting. We construct a set-valued map and show its value is a subset of the best-response policies, convex and upper hemi-continuous, which imply the existence of NE. We then introduce Robust Nash‑Iteration, and provide convergence guarantees. Finally, we connect average‑reward NE to discounted robust equilibria, showing approximation as the discount factor approaches one. Our studies provide comprehensive theoretical and algorithmic foundation for decision-making in complex, uncertain, and long-running multi-player environments.
@@ 1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments (shapley1953stochastic; filar2012competitive; littman1994markov). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the "Sim-to-Real" gap (mcmahan2024roping), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks (bukharin2023robust; ma2023decentralized). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
@@
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics (krnjaic2024scalable), communication networks (qu2020scalable), autonomous vehicle coordination (peake2020multi), financial markets (vadori2024towards), or even peer-to-peer energy trading (qiu2021multi). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate (hernandez2000zero). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium (filar2012competitive); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions (puterman1994markov; atia2021steady). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
@@ 1.1 Contributions
Solvability of the Robust Bellman Equation under Irreducible Setting: As a foundational step, we first revisit the single-agent distributionally robust average-reward MDPs with the irreducible assumption. Despite existing studies largely utilizing the robust Bellman equation (wang2023robust; wang2023model; wang2024robust; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample), its solvability is not proved, but rather implicitly assumed. We establish the well-posedness of the robust average-reward Bellman equation, study the structure of its solutions, and connect them with the optimal robust policies. This analysis provides structural insights and tools that are essential for solving multi-agent Markov games when long-term performance is of interest.
Existence and Algorithmic Identification of Robust Nash Equilibrium: We further introduce a formulation of induced distributionally robust MDPs to reduce the Nash Equilibrium identification in comprehensive multi-agent games to optimizing single-agent distributionally robust MDPs. Under the irreducibility setting, we utilize the solvability results we obtained to show a correspondence between the optimal robust policy and solutions of the Bellman equation, and prove the convexity and hemi-continuity of the best-response set, deriving the existence of a stationary Nash Equilibrium. We further study the more general weak communicating setting, where no such correspondence can be obtained. We hence instead study a proxy mapping which is a subset of the best-response set, and derive the NE based on it. These results address the fundamental open question, demonstrating that stable solutions exist in these complex, robust, and long-term horizon competitive environments.
Convergence Algorithm: We moreover propose a practical, iterative algorithm—a robust Nash iteration scheme—designed to compute the robust NE in the average-reward setting. We provide a rigorous proof of the algorithm’s convergence to the robust equilibrium policies, offering the first provably convergent method for solving average-reward DR-MGs.
Connection to Discounted Equilibria: We also aim to bridge the gap between DR-MGs under the average reward and the well-studied discounted-reward settings. We demonstrate the closeness between the robust average reward NE and robust discounted NEs as the discount factor approaches one. This connection not only provides deeper insight into the nature of the average reward equilibrium, but also suggests potential approximation methods based on the discounted framework.
@@ 1.2 Related Works
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, distributionally robust MDPs are first developed in (iyengar2005robust; nilim2004robustness) for both finite and discounted settings. The studies under the average reward setting are recently developed in (wang2023robust; wang2023model; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample; grand2023beyond; grand2023reducing; chatterjee2023solving; wang2024robust). While the robust Bellman equation is central to these methods, its solvability was often implicitly assumed without rigorous proof.
@@ 2 Preliminaries and Problem Formulation
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (littman1994markov), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\otimes_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}_{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the actions of all agents. At some state $s$, all agents act simultaneously by taking joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on stationary Nash Equilibrium (NE). More specifically, a stationary NE is a product policy $\pi^{\star}=(\pi^{\star}_{1},...,\pi^{\star}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{\star}$:
@@
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\otimes_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ (iyengar2005robust; nilim2004robustness; zhang2020robust; blanchet2023double; shi2024sample), where the uncertainty set is independently defined over all $(s,a)$ pairs.
After all agents take their actions $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}(\rho)$, and robust average reward $g^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}(\rho)$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{\star}$ such that for some initial state distribution $\rho$,
@@ 3 Hardness in Average Reward Nash Equilibrium Identification
Extending existing methods for discounted or finite-horizon robust MGs to the average-reward setting is prevented by two major challenges:
@@
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho).$ (4)
@@
To better illustrate the hardness of solving average reward DR-MGs, we first show the following impossibility result. The constructed DR-MG is essentially a multi-chain (puterman1994markov), indicating the necessity of the additional structural assumptions to solve average reward DR-MGs.
@@
There exists a two-player DR-MG without any stationary NE under average reward.
@@ 4 NE of Irreducible DR-MGs
As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs. In non-robust average reward MG studies, the NE can be proved by studying the induced stationary distribution (wikecek2015stationary). However, when considering robustness, the worst-case stationary distribution set becomes much more complicated. For example, it is non-convex (wang2022robust; zhang2024distributionally; ma2025rectified), whereas the convexity is necessary in non-robust proofs and hence the method cannot be adapted. We thus propose a direct approach by studying the best response policy.
@@
(Best Response Policy) In an average-reward DR-MG, fix agent $i$ and the joint policy $\pi_{-i}$ of all other agents. The set of best response policies of agent $i$ w.r.t. $\pi_{-i}$ is defined as
@@
$\pi^{\star}$ is a Nash Equilibrium if $\pi^{\star}_{i}\in\text{BR}_{i}(\pi^{\star}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{\star}$ is a fixed point of the BR mapping: $\pi^{\star}\in\text{BR}(\pi^{\star})$, where $\text{BR}(\pi)=\otimes_{i}\text{BR}_{i}(\pi^{\star}_{-i})$.
@@
Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) $f(x)$ is nonempty and convex $\forall x$; (2) $f$ is upper hemi-continuous (i.e., for any convergent sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$, it holds that $y\in f(x)$). Then there exists some $x^{\star}\in X$ such that $x^{\star}\in f(x^{\star})$.
@@
It is hence sufficient to fix an agent $i$ and study the best response policy set $\text{BR}_{i}(\cdot)$, which is a single-agent problem.
@@ 4.1 Induced Distributionally Robust MDPs
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, due to being induced, irreducibility, convexity, and compactness are all preserved. Namely, if the distributionally robust Markov game $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 1, then the induced distributionally robust MDP $\mathsf{M}_{i}(\pi_{-i})$ satisfies Assumption 1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
However, even the single-agent average reward robust RL is not fully studied. Specifically, the most essential tool, its robust Bellman equation, is not well studied: its solvability is not yet justified (an incorrect proof is given in (wang2023robust)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Although a huge body of existing work develops convergent algorithms, their convergence guaranties either rely on the wrong result in (wang2023robust), e.g., (wang2023model; Roch2025reduction), or bypass its solvability (xu2025finite; xu2025efficient; chatterjee2023solving; meggendorfer2025solving). We hence derive the solvability of the robust Bellman equation under Assumption 1 as follows.
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$; Moreover, for any solution $(g,V)$ to (7), it holds that $g=g^{\pi}_{\mathcal{P}},V=h^{\pi}_{\mathsf{P}_{w}}+c\textbf{e}$, for some constant $c$ and $\textbf{e}=(1,...,1).$
@@
Our results therefore prove the solvability of the robust Bellman equation, solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (7) can have non-unique solutions even up to constant vectors. Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward (puterman1990markov), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.

### v3

@@ Abstract
We study distributionally robust Markov games (DR‑MGs) with the average‑reward criterion, a framework for multi-agent decision-making under uncertainty over extended horizons. In average reward DR-MGs, agents aim to maximize their worst-case infinite-horizon average reward, to ensure satisfactory performance under environment uncertainties and opponent actions. We first establish a connection between the best-response policies and the optimal policies for the induced single‑agent problems. Under a standard irreducible assumption, we derive a correspondence between the optimal policies and the solutions of the robust Bellman equation, and derive the existence of stationary Nash Equilibrium (NE) based on these results. We further study DR-MGs under the weakly communicating setting, where We construct a set-valued map and show its value is a subset of the best-response policies, convex and upper hemi-continuous, and derive the existence of NE. We then explore algorithmic solutions, by first proposing a Robust Nash‑Iteration algorithm and providing convergence guarantees under some additional assumptions and a NE computing oracle. We further develop a temporal-difference based algorithm for DR-MGs, and provide convergence guarantees without any additional oracle or assumptions. Finally, we connect average‑reward robust NE to discounted ones, showing that the average reward robust NE can be approximated by the discounted ones under a large discount factor. Our studies provide a comprehensive theoretical and algorithmic foundation for decision-making in complex, uncertain, and long-running multi-player environments.
@@ 1 Introduction
A Markov Game (MG) or stochastic game provides a powerful mathematical framework for modeling sequential decision-making in competitive multi-agent environments (shapley1953stochastic; filar2012competitive; littman1994markov). Solving a MG involves finding an equilibrium policy profile for all players, where each player aims to maximize their own cumulative reward. However, a critical challenge in practice is the potential for a mismatch between the assumed MG model and the true underlying environment. This discrepancy, commonly called the Sim-to-Real gap (mcmahan2024roping), can arise from various sources, including environmental non-stationarity, inherent modeling errors, exogenous disturbances, or even adversarial attacks (bukharin2023robust; ma2023decentralized). Consequently, policies derived from a misspecified model can exhibit significantly degraded performance when deployed.
@@
In these formulations, the influence of future rewards diminishes exponentially, which often simplifies analysis as the results are less dependent on the long-term chain structure of the underlying game. However, despite the analytical elegance of these settings, policies learned under them can be suboptimal for multi-agent systems designed to operate over extended or indefinite time horizons, such as in warehouse robotics (krnjaic2024scalable), communication networks (qu2020scalable), autonomous vehicle coordination (peake2020multi), financial markets (vadori2024towards), or even peer-to-peer energy trading (qiu2021multi). In these scenarios, the long-term average reward is a more natural and meaningful performance measure. However, the study of DR-MGs under the average-reward criterion is nascent and fraught with complexities not present in the discounted or finite-horizon cases. The analysis of average-reward games hinges on the limiting behavior of the underlying stochastic process, making it markedly more intricate (hernandez2000zero). For instance, a general non-robust vanilla Markov game with average reward may not have stationary Nash Equilibrium (filar2012competitive); the direct correspondence between stationary policies and the limit points of state-action frequencies, a cornerstone of discounted analysis, breaks down in the average-reward setting even for single-agent MDPs, except under restrictive structural assumptions (puterman2014markov; atia2021steady). This difficulty is compounded in the multi-agent context, where strategic interactions are intertwined with the uncertain, complex dynamics.
@@ 1.1 Contributions
Existence of Robust Nash Equilibrium: We first study the solvability of average reward DR-MGs, i.e., the existence of robust NE. We first note that, finding the best response policy of an agent is equivelent to find the optimal policy of the single-agent induced robust MDP. We thus revisit the robust Bellman equation under average reward, and show that its solvability and the corresponding between its solutions and the optimal policies, under the irreducible assumption. Based on these results, we further prove the convexity and hemi-continuity of the best-response mapping, deriving the existence of a stationary Nash Equilibrium. We further study the more general weak communicating DR-MGs, where no such correspondence can be obtained due to their more complex structures. We instead study a proxy mapping which is a subset of the best-response mapping, and show the existence of NE based on it. These results address the fundamental open question, demonstrating that solutions exist in these complex, robust, and long-term horizon competitive environments.
Convergence Algorithm: We then look for concrete algorithms to find robust NE in average reward DR-MGs. Insipired by the standard value iteration approach, we first propose a robust Nash iteration algorithm. Under some additional assumptions on the game structure and an NE computing oracle, we prove the convergence to a robust NE of our algorithm. Moreover, to bypass these assumptions, we further develop a temporal-difference based method, which applies proximal descent on the temporal difference. Our algorithm does not require any NE oracle, and can be executed efficiently. We further develop its provably convergence guarantee without any additional assumption. Our algorithm designs and results hence provide the first comprehensive and algorithmic understanding on solving average-reward DR-MGs.
Connection to Discounted Equilibria: We further study the connection between the DR-MGs under average and discounted reward. We show that, the robust NE under the average reward can be effectively approximated by the robust NE under the discounted reward with some discount factor large enough. Such a connection enables us to solve the complicated and challenging average reward DR-MGs through the discounted ones, utilizing their element mathematical properties and extensively developed algorithms.
@@ 1.2 Related Works
Distributionally Robust MDPs. To address the performance degradation that occurs when a model is misspecified, distributionally robust MDPs are first developed in (iyengar2005robust; nilim2004robustness) for both finite and discounted settings. The studies under the average reward setting are recently developed in (wang2023robust; wang2023model; xu2025finite; xu2025efficient; roch2025finite; Roch2025reduction; chen2025sample; grand2023beyond; grand2023reducing; chatterjee2023solving; wang2024robust). However, these works are all developed for single-agent settings, and do not address the challenges we faced in multi-agent DR-MGs.
@@ 2 Preliminaries and Problem Formulation
Markov Games. Markov games (MGs) are the extension of MDPs under the multi-agent setting (littman1994markov), specified by $(\mathcal{N},\mathcal{S},\mathcal{A}=\times_{i\in\mathcal{N}}\mathcal{A}_{i},\{r_{i}\}_{i\in\mathcal{N}},\mathsf{P})$, where $\mathcal{N}=\{1,2,\dots,N\}$ is the set of all agents or players, $\mathcal{S}$ is the state space that all agents operate in, $\mathcal{A}_{i}$ denotes agent $i$’s action space. In a MG, both rewards and state transitions are determined by the joint actions of all agents. At some state $s$, all agents act simultaneously by taking a joint action $a=(a_{1},\dots,a_{N})\in\mathcal{A}$. Agent $i$ receives a reward of $r_{i}(s,a)$ for all $i\in\mathcal{N}$, and the next state is generated following the transition kernel $\mathsf{P}^{a}_{s}$.
In MGs, each agent $i$ has its own policy $\pi_{i}$, resulting in a joint (product) policy $\pi=(\pi_{1},...,\pi_{N})$. Each agent shares different goals as they have different reward functions to optimize their own value function/average reward based on its own reward and the joint policy: $V^{\pi}_{\mathsf{P},i}$ or $g^{\pi}_{\mathsf{P},i}$. Due to possible conflicting goals between agents, a MG aims to find some equilibria among all agents. In this work, we focus our study on stationary Nash Equilibrium (NE). More specifically, a stationary NE is a product policy $\pi^{\star}=(\pi^{\star}_{1},...,\pi^{\star}_{N})$, such that no single agent can improve its objective function by deviating from it, while the others stick to $\pi^{\star}$:
@@
Distributionally Robust Markov Games. Different from MGs, distributionally robust Markov games (DR-MGs) do not have a fixed transition kernel, but rather an uncertainty set $\mathcal{P}$ of transition kernels. We consider the standard $(s,a)$-rectangular uncertainty set $\mathcal{P}=\times_{(s,a)\in\mathcal{S}\times\mathcal{A}}\mathcal{P}^{a}_{s}$ (iyengar2005robust; nilim2004robustness; zhang2020robust; blanchet2023double; shi2024sample), where the uncertainty set is independently defined over all $(s,a)$ pairs.
After all agents take their joint action $a$ at state $s$, the next state is determined by an arbitrary kernel $\mathsf{P}^{a}_{s}\in\mathcal{P}^{a}_{s}$. In DR-MGs, agents adopt the principle of pessimism and optimize for the worst-case performance, which is defined as the robust value function $V^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}V^{\pi}_{\mathsf{P},i}(\rho)$, and robust average reward $g^{\pi}_{\mathcal{P},i}(\rho)\triangleq\min_{\mathsf{P}\in\mathcal{P}}g^{\pi}_{\mathsf{P},i}(\rho)$. The notion of Nash Equilibrium is similarly defined in terms of worst-case performance. In this paper, we focus on DR-MGs with average reward whose Nash Equilibrium is some product policy $\pi^{\star}$ such that for some initial state distribution $\rho$,
@@ 3 Hardness in Average Reward Nash Equilibrium Identification
Extending existing studies for discounted or finite-horizon robust MGs to the average-reward setting is prevented by two major challenges:
@@
$\max_{\pi_{1}}\min_{\pi_{2}}\min_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho)\neq\min_{\pi_{2}}\max_{\pi_{1}}\max_{\mathsf{P}\in\mathcal{P}}V^{(\pi_{1},\pi_{2})}_{\mathsf{P},r}(\rho).$
@@
To better illustrate the hardness of solving average reward DR-MGs, we first show the following impossibility result. The constructed DR-MG is essentially a multi-chain (puterman2014markov), indicating the necessity of the additional structural assumptions to solve average reward DR-MGs.
@@
There exists a DR-MG without any stationary robust NE under average reward.
@@ 4 NE of Irreducible DR-MGs
As discussed in Section 3, neither the limit method nor the Min-Max duality is applicable in average-reward DR-MGs. In non-robust average reward MG studies, the NE can be proved by studying the induced stationary/occupancy distribution (wikecek2015stationary). However, when considering robustness, the worst-case stationary distribution set becomes much more complicated. For example, it is non-convex (wang2022robust; zhang2024distributionally; ma2025rectified), whereas the convexity is necessary in non-robust proofs and hence the method cannot be adapted. We thus propose a direct approach by studying the best response mapping.
@@
(Best Response mapping) In an average-reward DR-MG, fix agent $i$ and the joint policy $\pi_{-i}$ of all other agents. The best response mapping of agent $i$ w.r.t. $\pi_{-i}$ is a set-valued mapping defined as
@@
$\pi^{\star}$ is a Nash Equilibrium if and only if $\pi^{\star}_{i}\in\text{BR}_{i}(\pi^{\star}_{-i})$, $\forall i\in\mathcal{N}$; or equivalently, if $\pi^{\star}$ is a fixed point of the BR mapping: $\pi^{\star}\in\text{BR}(\pi^{\star})$, where $\text{BR}(\pi)=\times_{i}\text{BR}_{i}(\pi^{\star}_{-i})$.
@@
Let $X$ be a compact convex subset of $\mathbb{R}^{n}$ and let $f:X\to 2^{X}$ be a set-valued function for which, (1) $f(x)$ is nonempty and convex, $\forall x$; (2) $f$ is upper hemi-continuous (i.e., for any convergent sequences $\{x_{n}\}\to x$ and $\{y_{n}\}\to y$ such that $y_{n}\in f(x_{n})$, it holds that $y\in f(x)$). Then there exists some $x^{\star}\in X$ such that $x^{\star}\in f(x^{\star})$.
@@ 4.1 Induced Distributionally Robust MDPs
Clearly, $(\mathcal{S},\mathcal{A}_{i},\mathcal{P}^{\pi_{-i}},r^{\pi_{-i}})$ is a single-agent distributionally robust MDP for agent $i$. The uncertainty set $\mathcal{P}^{\pi_{-i}}$ is $(s,a_{i})$-rectangular, due to the $(s,a)$-rectangularity of the underlying uncertainty set $\mathcal{P}$. Moreover, due to being induced, irreducibility, convexity, and compactness are all preserved. Namely, if the DR-MG $\mathcal{MG}$ satisfies the multi-agent extension of Assumption 1, then the induced distributionally robust MDP $\mathsf{M}_{i}(\pi_{-i})$ satisfies Assumption 1. Thus all of the previous results for the single-agent setting are still valid under the induced distributionally robust MDP for a fixed policy $\pi_{-i}$, based on which we can show the existence of a NE.
However, even the single-agent average reward robust RL is not fully studied. Specifically, the most essential tool, its robust Bellman equation, is not well studied: its solvability or characterizations of solutions are not yet justified (an incorrect proof is given in (wang2023robust)), therefore making the connection between optimal policies and the robust Bellman equation unclear. Although a huge body of existing work develops convergent algorithms, their convergence guaranties either rely on the wrong result in (wang2023robust), e.g., (wang2023model; Roch2025reduction), or bypass its solvability (xu2025finite; xu2025efficient; chatterjee2023solving; meggendorfer2025solving). We hence derive the solvability of the robust Bellman equation under Assumption 1 as follows.
@@
where $\sigma_{\mathcal{P}^{a}_{s}}(V)\triangleq\min_{P\in\mathcal{P}^{a}_{s}}PV$; The uniqueness is in the sense that, for any solution $(g,V)$ to (6), it holds that $g=g^{\pi}_{\mathcal{P}},V=h^{\pi}_{\mathsf{P}_{w}}+c\textbf{e}$, for some constant $c$ and $\textbf{e}=(1,...,1).$
@@
Our results therefore prove the solvability of the robust Bellman equation, especially (6), solidifying the connection between the robust average reward and its solutions. We note that the robust Bellman equation can have a more complicated structure than the non-robust one: any worst-case kernel $\mathsf{P}_{w}$ will result in a solution, hence (6) may have non-unique solutions even up to constant vectors (see discussion in wang2023model). Our proof techniques are also different from the non-robust ones, which rely on the Laurent series of the non-robust average reward (puterman2014markov), whereas we directly show the constructed pair $(g^{\pi}_{\mathsf{P}_{w}},h^{\pi}_{\mathsf{P}_{w}})$ is a solution.